OpenStudy (sammieell):

given QRS congruent to TUV, QS=3v+2 and TV=7v-6 find the length of QS and TV

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) 3v+2=7v-6 solve for v

jigglypuff314 (jigglypuff314):

3v+2=7v-6 solve for v try subtracting 3v from both sides then adding 6 to both sides then dividing both sides by 4 do you know how to do this?

OpenStudy (sammieell):

no can you explain it more

jigglypuff314 (jigglypuff314):

by saying that QRS is congruent to TUV that means QS = TV so 3v+2 = 7v-6 follow?

OpenStudy (sammieell):

yes

jigglypuff314 (jigglypuff314):

so then solve for v from there

jigglypuff314 (jigglypuff314):

3v+2 = 7v-6 -3v -3v so 2 = 4v - 6 then +6 +6

OpenStudy (sammieell):

i understand now

jigglypuff314 (jigglypuff314):

good :) once you get v you can then plug it back into QS = 3v + 2 = and TV = 7v - 6 =

OpenStudy (sammieell):

they both equal 8

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (sammieell):

i need help with another problem though

jigglypuff314 (jigglypuff314):

sure, ask away :)

OpenStudy (sammieell):

what other information do you need in order to prove the triangles congruent using SAS Congruence Postulate

jigglypuff314 (jigglypuff314):

AC = AC because they are the same side BC = CD is given then you would need the angle between them to be congruent if you want a SAS

OpenStudy (sammieell):

how do i get that

jigglypuff314 (jigglypuff314):

hmmm?

jigglypuff314 (jigglypuff314):

http://prntscr.com/53tbg3 name those two angles

OpenStudy (sammieell):

BD ?

jigglypuff314 (jigglypuff314):

angle ACB = angle ACD

OpenStudy (sammieell):

oh ok

OpenStudy (sammieell):

thank you

jigglypuff314 (jigglypuff314):

Glad I could help :)